OpenStudy (jimbo69):

Solve the given proportion. 1/9= (x+3)/2

Directrix (directrix):

1/9= (x+3)/2 Use the Means-Extremes Property

Directrix (directrix):

1*2 = 9(x + 3) @Jimbo69 Would you post the next step?

OpenStudy (vijeya3):

  If this helpful,fan and medal me.Tag me whenever you need help..:)